Show all attempts Jar Jar forms an aggressive growth portfolio by investing 20% of his savings in Ford stock, 23% in Toyota sPlease show the steps to get the right answers. Thank you!
Answer #1

Solution:-

Calculation of portfolio beta:

The beta of a portfolio is simply the weighted average of betas of individual components of the portfolio. As a simple example, if a portfolio has two assets in equal proportion with asset A having a beta of 1.5 and asset B having a beta of 0.5, the portfolio beta will be as follows:

Portfolio beta (for above example)= Beta A*50% + Beta B*50%= 1.5*50% + 0.5*50%= 1

Thus, we can summarise the formula for calculating beta as follows:

Beta (portfolio of n assets)= (Beta Asset 1 * weight Asset 1) + (Beta Asset 2 * weight Asset 2) + (Beta Asset 3 * weight Asset 3) +...………..+ (Beta Asset n * weight Asset n)

Now, as given in the question we have the following information:

Beta (Ford)= 1.18 ; Weight (Ford)= 20%

Beta (Toyota)= 1.41 ; Weight (Toyota)= 23%

Beta (Kia)= 1.81 ; Weight (Toyota)= 24%

Beta (Market index)= 1 ; Weight (Toyota)= 10%

Beta (Risk free bond)= 0 (Betas of risk free assets are zero) ; Weight (Toyota)= 23%

Using the above information and beta formula as shown above, we calculate beta portfolio as follows:

Beta (Portfolio)= (1.18*20%) + (1.41*23%) + (1.81*24%) + (1*10%) + (0*23%)

Beta (Portfolio)= 1.09

Calculation of portfolio expected return:

As per CAPM, the required rate of return is calculated as follows:

Expected Return= Risk free rate + Beta*Market risk premium

Risk-free rate (Equal to return of risk free bond)= 7%

Market risk premium= Expected return market index - risk free rate= 10%-7%= 3%

Portfolio beta= 1.09

Thus, the expected returns (ER) for portfolio is calculated using CAPM as follows:

ER (Portfolio)= 7% + 1.09*3% = 10.27%